# AgileRate: Bringing Adaptivity and Robustness to DeFi Lending Markets

### v3

@@ 1. Introduction
The following sections summarize our results and set the broader research context in which we make these contributions.
@@ 3.1. Market participants and pool mechanics
Borrowers Borrowers secure loans by providing collateral in the form of an asset $\mathcal{A}_{c}$, which backs the loan they take in a different asset, $\mathcal{A}_{l}$, supplied by lenders through a pooled system. The loan amount for borrower $i$ at time $t$ is represented by $B_{t}(i)$, while the total amount borrowed by all borrowers is $B_{t}$. The collateral posted by borrower $i$ and the total collateral in $\mathcal{A}_{c}$ are denoted by $C_{t}(i)$ and $C_{t}$, respectively. To open a borrowing position, a borrower must satisfy the condition $\frac{B_{t}(i)}{C_{t}(i)p_{t}}<c_{t}<1$, where $p_{t}$ is the price of $\mathcal{A}_{c}$ relative to $\mathcal{A}_{l}$ at time $t$, and $c_{t}$ is the protocol-defined collateral factor at time $t$. The protocol also defines an interest rate, $r_{t}$, that applies to all open borrowing positions in each timeslot, causing the total borrowed amount $B_{t}$ to increase over time as interest accrues.
Lenders Lenders deposit asset $\mathcal{A}_{l}$ to earn interest on the amounts borrowed by the borrowers. In this analysis, we focus on a lending protocol where all accrued interest goes directly to the lenders. Unlike many existing platforms, which reserve a portion of the interest for a fund to cover potential defaults and distribute some as rewards to token holders (aaveWP, [n. d.]; compoundWP, [n. d.]), the platform we study has no such reserves. As a result, any borrower default directly reduces the deposits of lenders, transferring the risk of default from the protocol to the lenders. Let $L_{t}(i)$ be the deposit of lender $i$ in $\mathcal{A}_{l}$, and let $L_{t}$ be the total deposits. The utilization rate, $U_{t}=\frac{B_{t}}{L_{t}}$, expresses the fraction of deposits currently borrowed. Lenders earn interest at rate $r_{t}$ on the portion of their deposits that are utilized. In the event of a default, each lender’s loss is proportional to their original deposit relative to the total pool.
Liquidators To protect the system against falling collateral asset prices, a liquidation mechanism is implemented, similar to those found on platforms like Aave and Compound. When the loan-to-value (LTV) ratio, which we previously defined as the fraction of debt to collateral, exceeds the liquidation threshold $LT_{t}<1$, the borrowing position can be liquidated. This allows a third-party liquidator to interact with the smart contract, repay part of the borrower’s debt in $\mathcal{A}_{l}$, and claim an equivalent portion of the borrower’s collateral in $\mathcal{A}_{c}$, along with an incentive fee, $LI_{t}$, to reward the liquidator. Liquidation is possible until the LTV ratio drops back below the threshold $LT_{t}$.
@@ 3.4.1. Demand-supply balancing
The expectation is taken over the protocol’s randomness, as well as the noise present in the users’ demand and supply.  [^1] A more pragmatic version of demand-supply balancing that maximizes the revenue of the protocol is formally defined in Appendix [A.1] and Appendix [A.2]
@@ 3.4.3. Liquidation and default risk control
Pool defaults occur when a borrower’s debt exceeds the value of their collateral due to a price drop, making the pool unable to recover the full loan amount. Liquidation occurs when a borrower, maintaining the maximum loan-to-value ratio, faces a price drop that forces the protocol to liquidate part of their collateral to bring the ratio back below the liquidation threshold. These risks are formally defined in Appendix [A.5], where the exact conditions for pool default and liquidation are detailed.
@@ 4.1. Interest rate controller
Compute the gain vector for the supply and demand observation :
Update the parameter estimate:
Update the covariance matrix:
end for
Optimizing interest rate for desired utilization Given the demand and supply models presented in Equations 11 and 12, when the noise is small, the utilization rate can be approximated with:
@@
The objective of our protocol is to set the interest rate $r_{t}$ such that the utilization remains close to a desired utilization $U^{*}$. The optimal interest rate $r^{*}_{\text{utl}}$ that achieves this desired utilization can be derived as follows:  [^2] We develop similar results for the objective of maximizing revenue instead of optimizing rate deviation. See Appendix [A.2]
@@ 5.1. Learning-based interest rate controller
Learning-based interest rate controllers estimate demand and supply while incorporating prior beliefs to set optimal rates. As adaptivity increases, these controllers become more vulnerable to manipulation, quickly updating beliefs based on a few manipulated samples. To quantify this impact, we analyze an adaptive learning-based algorithm, $\mathtt{LC}$, which alternates between exploration and exploitation phases. During exploration, the controller detects discrepancies between observed and expected demand and supply, testing various rates to update its beliefs. Once stabilized, the controller enters exploitation, setting rates based on the learned functions. This cycle repeats as new changes are detected. We model $\mathtt{LC}$ as an auction-like mechanism. In exploration, borrowers and lenders report their minimum (or maximum) acceptable rates and quantities. $\mathtt{LC}$ infers the demand and supply curves from this data. In exploitation, it sets a single rate $r_{t}$, allowing lenders with bids $r_{t}$ or lower to deposit, and borrowers with bids $r_{t}$ or higher to receive loans. This mimics a uniform price auction, where participants’ reported bids influence the interest rate, maximizing utility based on private valuations. Participant utility depends on whether the rate falls within their acceptable range and the portion of their supply or demand allocated to the protocol or external markets. More details on this auction implementation are provided in Appendix [A.3].
@@ 5.2. Static interest rate curves
Let a strategic borrower $\mathcal{A}^{b}$, with external rate $r^{b}$, control $\delta_{b}$ of the total demand $B$, and a strategic lender $\mathcal{A}^{l}$, with external rate $r^{l}$, control $\delta_{l}$ of the total supply $L$. Both aim to maximize their utility by adjusting how much they participate in the pool. The full utility functions for strategic borrowers and lenders are provided in Appendix [A.4].
@@
Given the static curve described in 10, with fixed truthful demand $(1-\delta_{b})B$ and fixed truthful supply $(1-\delta_{l})L$, and the remaining demand $\delta_{b}B$ and supply $\delta_{l}L$ controlled by strategic borrowers and lenders, the Adversarial Impact is bounded by:
@@ 5.3.1. Ecosystem modification
In Theorem 5.2, we have shown that the interest rate of a lending pool that follows an adaptive algorithm is susceptible to adversarial manipulation via strategic withholding. We now see if there are ways of mitigating this kind of a manipulation.
Note that the core of the problem stems from the fact that the demand (or supply) of liquidity to the pool reacts in a delayed fashion to deviation of the demand (or supply) from the norm. By the “norm”, we mean the demand (or supply) that borrowers (or lenders) have for the asset if everyone was honest. This delay is captured by the parameter of demand (or supply) elasticity. A more elastic demand implies a more reactive market. This, in turn, implies a higher chance of the demand seen by the lending pool to be a true representation of market demand. The question of mitigating adversarial effects on demand and supply boils down to the question of increasing supply and demand elasticities of the underlying asset with respect to the interest rates. This increase can be achieved by the following strategies.
Naively, the elasticity of a pool can be boosted by exposing the pool to as much alternative trading volume as possible. The presence of multiple competing lending pools would increase elasticity, but is rarely in the hands of the lending pool designer. Additionally, the designer can replicate the pool across different blockchain platforms including Layer-1’s and Layer-2’s. The presence of fast and robust bridges across all of these platforms would help unify liquidity and encourage arbitrage and hence increase pool elasticity.
Alternatively, the presence of a deep derivatives market for interest rates of specific pools in a lending protocol will also help increase demand/supply elasticity. A derivatives market enables speculators and arbitrageurs to take note of discrepancies in what the futures markets expect the interest rates to be and what it actually is due to adversarial manipulation. In traditional finance, for example, the use of credit derivatives has been observed to be associated with an increased supply of bank credit (Hirtle, 2008). On the demand side, the use of credit derivatives such as interest rate swaps allow more sophisticated users to face a variable interest rate while less sophisticated users can fix their desired interest rate. This allows a user more capable of managing an adaptive interest rate to take on the risk entailed by such a rate. This user would also more sensitive in assessing whether the current interest rate offered by a pool is reasonable, and takes an action to borrow more if the rate is too low. This, in turn, would increase the demand elasticity of the interest rate market. Markets offering yield tokenization derivatives such as Pendle (Finance, 2022), or fixed rate lending/borrowing such as IPOR (Camas et al., 2021) effectively allow users in DeFi to enter interest rate swaps. Other protocols such as Timeswap (Ngo, 2022) effectively offer a futures market on interest rates. In principle, a sophisticated trader should be able to use the prices of swaps on these external markets and assess whether the adaptive interest rate offered on a lending pool is fair. A transparent valuation model for all such protocols is an interesting open problem and would help boost elasticity and hence protocol robustness for all lending pools, but is beyond the scope of this paper. We leave the pricing of such derivatives to future work.
@@ 5.3.2. Algorithmic mitigation
where $\phi(.)$ is the M-estimate function. The M-estimate function is a redescending function, meaning its derivative decreases to zero as the residual $e$ becomes very large. Refer to Appendix A.6 for the description of the M-estimate function we adopt. Zou et al. show that minimizing the robust loss function in Equation 17 results in a recursive algorithm similar to the standard RLS with forgetting factor, except that the gain is updated as: $\mathbf{K}_{t}=\frac{q(e_{t})\mathbf{P}_{t-1}\mathbf{x}_{t}}{\rho+q(e_{t})\mathbf{x}_{t}^{T}\mathbf{P}_{t-1}\mathbf{x}_{t}},$ where $e_{t}\coloneqq(y_{t}-\mathbf{x}_{t}^{T}\hat{\bm{\theta}}_{t-1})^{2}$ and $q(e)\coloneqq\frac{d\phi(e)}{de}/e$.
@@ 6.1. Empirical validation of demand and supply models:
We applied our RLS-based algorithm to demand and supply data from Compound market pools, analyzing the relationship between interest rates and the demand and supply curves in 3-hour intervals. Specifically, we fitted the data (Feb 2024 - Feb 2025, RLS with $\rho=0.95$) to our demand and supply models (Equations 12 and 11) and estimated the parameters $\mathtt{a}_{t}^{b}$, $\mathtt{b}_{t}^{b}$, $\mathtt{a}_{t}^{l}$, and $\mathtt{b}_{t}^{l}$ over time. To evaluate the accuracy of these parameter estimates, we predicted demand and supply for the next timeslot and summarized the prediction errors in Table 1. The relatively low error across the three main pools demonstrates the effectiveness of our model and methodology. More details about our experiment are provided in the Appendix B.1.
@@ 6.2. Interest rate controller
To evaluate our interest rate controller, we analyze two demand-supply models: one based on real-world Compound data (Section 6.1) and another using synthetic parameter trajectories. The latter follows a Gaussian random walk, where each parameter ($\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$) evolves every $T$ steps as $\mathtt{a}^{b}_{t+T}=\mathtt{a}_{t}^{b}+\epsilon_{t}$, with $\epsilon_{t}\sim\mathcal{N}(0,\sigma_{\text{trns}}\cdot|\mathtt{a}_{t}^{b}|)$. Demand and supply are then generated using Equations 11 and 12, incorporating normally distributed noise ($\sigma_{\text{noise}}$).
@@
Utilization error We evaluate our RLS-based controller against Compound’s static interest rate curves tuned for optimal utilization of $0.8$. Using Compound’s data (Feb 2024–Feb 2025), we infer $\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$ to simulate user behavior. Our controller archives significantly lower utilization mean squared error (MSE) across all pools (Table 3). We use $\rho=0.85$. To assess robustness, we test synthetic data. As $\sigma_{\text{trns}}$ increases, inducing greater demand-supply fluctuations, our controller incurs more errors but consistently outperforms the Aave/Compound static approach (Figure 2(b)). For details, see Appendix B.2.
@@ 7. Conclusion
Appendix A Supplementary explanations about the model and the results
A.1. Alternate objective for demand/supply balancing

### v4

@@ 1. Introduction
The following section summarizes our results. We set the broader research context in which we make these contributions in Appendix [B.1].
@@ 2.1. Market participants and pool mechanics
Borrowers Borrowers secure loans by providing collateral in asset $\mathcal{A}_{c}$ to borrow asset $\mathcal{A}_{l}$ from a pooled system. The loan amount for borrower $i$ at time $t$ is $B_{t}(i)$, with total borrowings $B_{t}$, and their collateral is $C_{t}(i)$, with total collateral $C_{t}$. To open a borrowing position, borrowers must satisfy $\frac{B_{t}(i)}{C_{t}(i)p_{t}}<c_{t}<1$, where $p_{t}$ is the price of $\mathcal{A}_{c}$ relative to $\mathcal{A}_{l}$, and $c_{t}$ is the protocol-defined collateral factor. The protocol applies a global interest rate $r_{t}$ to all borrowings, increasing $B_{t}$ over time.
Lenders Lenders deposit $\mathcal{A}_{l}$ to earn interest from borrowers. Unlike many platforms that reserve interest for risk funds and token rewards (aaveWP, [n. d.]; compoundWP, [n. d.]), this protocol distributes all accrued interest to lenders, making them directly bear default risks. Let $L_{t}(i)$ be lender $i$’s deposit and $L_{t}$ be the total deposits, with utilization rate $U_{t}=\frac{B_{t}}{L_{t}}$. Lenders earn interest at $r_{t}$ on borrowed funds, and in case of default, losses are proportional to their deposits.
Liquidators To prevent risks from falling collateral prices, a liquidation mechanism allows third-party liquidators to intervene when a borrower’s loan-to-value (LTV) ratio exceeds the liquidation threshold $LT_{t}<1$. A liquidator repays part of the borrower’s debt in $\mathcal{A}_{l}$, claims a corresponding amount of $\mathcal{A}_{c}$, and earns an incentive fee $LI_{t}$. Liquidations continue until the LTV ratio falls below the liquidation threshold $LT_{t}$.
@@ 2.4. Protocol objectives
We define the protocol objectives in a three-fold manner. The explicit mathematical definitions we use for these objectives are given in Appendix [B.2].
@@ 2.4.1. Demand-supply balancing
In decentralized finance (DeFi) platforms, maintaining an optimal balance between supply and demand is crucial, with the key metric being proximity to the target utilization rate, $U^{*}$ (Group, 2025a, b). Low utilization implies inefficient use of deposits, while high utilization restricts withdrawals. To evaluate a protocol $\mathcal{P}$’s ability to maintain $U^{*}$, we define the optimal interest rate $r^{*}_{\text{utl}}$ as the rate minimizing deviation from $U^{*}$, given the utilization function $U=f(r;\mathcal{B}_{0})/g(r_{\text{eff}};\mathcal{L}_{0})$ where $r_{\text{eff}}=U\cdot r$. We introduce Rate Deviation, denoted by $\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*})$, as the expected deviation of the protocol’s interest rate from $r^{*}_{\text{utl}}$, considering randomness in protocol execution and noise in user demand and supply. A more revenue-centric version of this balancing mechanism is defined in Appendix [B.3] and Appendix [B.4].
@@ 2.4.2. Adversarial robustness
Strategic borrowers and lenders can manipulate a protocol’s interest rate by simulating demand and supply responses to maximize personal utility. We analyze the extent of such manipulation by considering a protocol $\mathcal{P}$ with truthful users following demand and supply curves $f(r;\mathcal{B}_{t})$ and $g(r_{\text{eff}};\mathcal{L}_{t})$, alongside a strategic lender $\mathcal{A}^{l}$ and a strategic borrower $\mathcal{A}^{b}$, who control a fraction $\delta_{l}$ and $\delta_{b}$ of maximum supply and demand, respectively, and have external rates $r^{l}$ and $r^{b}$. Defining $r_{\text{truthful}}$ and $r_{\text{strategic}}$ as the steady-state interest rates with and without strategic behavior, we introduce Adversarial Impact Deviation, denoted by $AI^{\mathcal{P}}(\delta_{b},\delta_{l})$, as the supremum of the expected deviation between these rates over all external rate choices of strategic users.
@@ 2.4.3. Liquidation and default risk control
In peer-to-pool-to-peer lending, the primary risks are borrower defaults and liquidations, requiring dynamic adjustments of parameters such as the collateral factor $c_{t}$, liquidation threshold $LT_{t}$, and liquidation incentive $LI_{t}$ based on price volatility. Defaults occur when a borrower’s debt exceeds collateral value due to price drops, while liquidations happen when a price decline forces the protocol to sell collateral to restore safe loan-to-value ratios. These risks are formally defined in Appendix [B.7]. The protocol’s ability to manage these risks can be quantified using metrics such as the expected value or 95th percentile of pool defaults and liquidations, computed with respect to the price distribution.
@@ 3.1. Interest rate controller
Optimizing interest rate for desired utilization Given the demand and supply models presented in Equations 8 and 9, when the noise is small, the utilization rate can be approximated with:
@@
The objective of our protocol is to set the interest rate $r_{t}$ such that the utilization remains close to a desired utilization $U^{*}$. The optimal interest rate $r^{*}_{\text{utl}}$ that achieves this desired utilization can be derived as follows:  [^1] We develop similar results for the objective of maximizing revenue instead of optimizing rate deviation. See Appendix [B.4]
@@ 4.1. Learning-based interest rate controller
Learning-based interest rate controllers estimate demand and supply while incorporating prior beliefs to set optimal rates. As adaptivity increases, these controllers become more vulnerable to manipulation, quickly updating beliefs based on a few manipulated samples. To quantify this impact, we analyze an adaptive learning-based algorithm, $\mathtt{LC}$, which alternates between exploration and exploitation phases. During exploration, the controller detects discrepancies between observed and expected demand and supply, testing various rates to update its beliefs. Once stabilized, the controller enters exploitation, setting rates based on the learned functions. This cycle repeats as new changes are detected. We model $\mathtt{LC}$ as an auction-like mechanism. In exploration, borrowers and lenders report their minimum (or maximum) acceptable rates and quantities. $\mathtt{LC}$ infers the demand and supply curves from this data. In exploitation, it sets a single rate $r_{t}$, allowing lenders with bids $r_{t}$ or lower to deposit, and borrowers with bids $r_{t}$ or higher to receive loans. This mimics a uniform price auction, where participants’ reported bids influence the interest rate, maximizing utility based on private valuations. Participant utility depends on whether the rate falls within their acceptable range and the portion of their supply or demand allocated to the protocol or external markets. More details on this auction implementation are provided in Appendix [B.5].
@@ 4.2. Static interest rate curves
Let a strategic borrower $\mathcal{A}^{b}$, with external rate $r^{b}$, control $\delta_{b}$ of the total demand $B$, and a strategic lender $\mathcal{A}^{l}$, with external rate $r^{l}$, control $\delta_{l}$ of the total supply $L$. Both aim to maximize their utility by adjusting how much they participate in the pool. The full utility functions for strategic borrowers and lenders are provided in Appendix [B.6].
@@
Given the static curve described in 7, with fixed truthful demand $(1-\delta_{b})B$ and fixed truthful supply $(1-\delta_{l})L$, and the remaining demand $\delta_{b}B$ and supply $\delta_{l}L$ controlled by strategic borrowers and lenders, the Adversarial Impact is bounded by:
@@ 4.3.1. Ecosystem modification
In Theorem 4.2, we showed that an adaptive interest rate in lending pools is vulnerable to adversarial manipulation via strategic withholding. The core issue arises from the delayed response of liquidity demand (or supply) to deviations from the norm, where elasticity plays a crucial role in mitigating such manipulation. Higher elasticity ensures that observed demand better reflects true market demand, reducing manipulation risks. To enhance elasticity, lending pool designers can increase exposure to alternative trading venues, deploy pools across multiple blockchain platforms with robust cross-chain bridges, and foster arbitrage opportunities. Additionally, deep derivatives markets for interest rates, such as those enabled by yield tokenization (e.g., Pendle (Finance, 2022)), fixed-rate lending (e.g., IPOR (Camas et al., 2021)), and futures markets (e.g., Timeswap (Ngo, 2022)), can help sophisticated traders assess fair interest rates and respond accordingly. By allowing speculators and arbitrageurs to act on discrepancies between expected and actual rates, these mechanisms increase demand and supply elasticity, making lending protocols more resilient. A transparent valuation model for such derivatives remains an open problem and is left for future research.
@@ 4.3.2. Algorithmic mitigation
where $\phi(.)$ is the M-estimate function. The M-estimate function is a redescending function, meaning its derivative decreases to zero as the residual $e$ becomes very large. Refer to Appendix B.8 for the description of the M-estimate function we adopt. Zou et al. show that minimizing the robust loss function in Equation 14 results in a recursive algorithm similar to the standard RLS with forgetting factor, except that the gain is updated as: $\mathbf{K}_{t}=\frac{q(e_{t})\mathbf{P}_{t-1}\mathbf{x}_{t}}{\rho+q(e_{t})\mathbf{x}_{t}^{T}\mathbf{P}_{t-1}\mathbf{x}_{t}},$ where $e_{t}\coloneqq(y_{t}-\mathbf{x}_{t}^{T}\hat{\bm{\theta}}_{t-1})^{2}$ and $q(e)\coloneqq\frac{d\phi(e)}{de}/e$.
@@ 5.1. Empirical validation of demand and supply models:
We applied our RLS-based algorithm to demand and supply data from Compound market pools, analyzing the relationship between interest rates and the demand and supply curves in 3-hour intervals. Specifically, we fitted the data (Feb 2024 - Feb 2025, RLS with $\rho=0.95$) to our demand and supply models (Equations 9 and 8) and estimated the parameters $\mathtt{a}_{t}^{b}$, $\mathtt{b}_{t}^{b}$, $\mathtt{a}_{t}^{l}$, and $\mathtt{b}_{t}^{l}$ over time. To evaluate the accuracy of these parameter estimates, we predicted demand and supply for the next timeslot and summarized the prediction errors in Table 1. The relatively low error across the three main pools demonstrates the effectiveness of our model and methodology. More details about our experiment are provided in the Appendix C.1.
@@ 5.2. Interest rate controller
To evaluate our interest rate controller, we analyze two demand-supply models: one based on real-world Compound data (Section 5.1) and another using synthetic parameter trajectories. The latter follows a Gaussian random walk, where each parameter ($\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$) evolves every $T$ steps as $\mathtt{a}^{b}_{t+T}=\mathtt{a}_{t}^{b}+\epsilon_{t}$, with $\epsilon_{t}\sim\mathcal{N}(0,\sigma_{\text{trns}}\cdot|\mathtt{a}_{t}^{b}|)$. Demand and supply are then generated using Equations 8 and 9, incorporating normally distributed noise ($\sigma_{\text{noise}}$).
@@
Utilization error We evaluate our RLS-based controller against Compound’s static interest rate curves tuned for optimal utilization of $0.8$. Using Compound’s data (Feb 2024–Feb 2025), we infer $\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$ to simulate user behavior. Our controller archives significantly lower utilization mean squared error (MSE) across all pools (Table 3). We use $\rho=0.85$. To assess robustness, we test synthetic data. As $\sigma_{\text{trns}}$ increases, inducing greater demand-supply fluctuations, our controller incurs more errors but consistently outperforms the Aave/Compound static approach (Figure 2(b)). For details, see Appendix C.2.
@@ 6. Conclusion
7. Acknowledgment
This work was supported by the National Science Foundation via grants CCF-1705007, CNS-2325477, the Army Research Office via grant W911NF2310147, a grant from C3.AI and a gift from XinFin Private Limited.
Appendix A Details for our algorithms
In this section, we detail the algorithms we use in Section 3.1.
Compute the gain vector for the supply and demand observation :
Update the parameter estimate:
Update the covariance matrix:
end for
end for
Appendix B Supplementary explanations about the model and the results
B.1. Related work
B.2. Explicit definition of various objectives
B.2.1. Demand-supply balancing
The expectation is taken over the protocol’s randomness, as well as the noise present in the users’ demand and supply.  [^2] A more pragmatic version of demand-supply balancing that maximizes the revenue of the protocol is formally defined in Appendix [B.3] and Appendix [B.4]
B.2.2. Adversarial robustness
B.2.3. Liquidation and default risk control
Pool defaults occur when a borrower’s debt exceeds the value of their collateral due to a price drop, making the pool unable to recover the full loan amount. Liquidation occurs when a borrower, maintaining the maximum loan-to-value ratio, faces a price drop that forces the protocol to liquidate part of their collateral to bring the ratio back below the liquidation threshold. These risks are formally defined in Appendix [B.7], where the exact conditions for pool default and liquidation are detailed.
B.3. Alternate objective for demand/supply balancing